OpenStudy (anonymous):

I need to find an equation that gives me a discriminant of 0 anybody have one?

OpenStudy (anonymous):

Welcome to openstudy.

OpenStudy (anonymous):

Do you know what a discriminant is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so what does 4^2 - 4(2)(2)

OpenStudy (anonymous):

oh we turn stuff like \[3x ^{2}-8x+12=0 \] into discriminant s using the quadratic equation

OpenStudy (anonymous):

HINT: It equals 0

OpenStudy (anonymous):

Yuh bruh

OpenStudy (anonymous):

so how would you write that equation

OpenStudy (anonymous):

4^2-4(2)(2) where b^2-4ac in ax^2+bx+c

OpenStudy (loser66):

You have the formula to find discriminant, it is \(b^2-4ac\) and you need it =0, so just let it =0, you get \(b^2-4ac=0 \) then \(b^2=4ac\), pick whatever you like to make them equal.

OpenStudy (anonymous):

so \[2x^{2}+4x+2\]?

OpenStudy (anonymous):

you mean a quadratic equation? any perfect square set equal to zero will for example \(z^2=0\) or \((x+3)^2=0\) or \((2x-1)^2=0\)

OpenStudy (anonymous):

i need it in \[ax ^{2}+bx+c\]

OpenStudy (anonymous):

ok then for example \[(x+3)^2=0\\ x^2+6x+9=0\]

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

in other words, don't think so hard about how to make \[b^2-4ac=0\] just square something

OpenStudy (anonymous):

im doing a semester exam review

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

oh little joy!

OpenStudy (anonymous):

Yeah... and i have to make the review myself so.... even better right?

OpenStudy (anonymous):

i think the final episode of gracepoint is on tonight, sound more interesting

OpenStudy (anonymous):

lol :D

OpenStudy (anonymous):

good luck with your exam why do you have to make your own review?

OpenStudy (anonymous):

IDK..... I guess so we can study the material while we make the review

OpenStudy (anonymous):

sounds like when you have to make your own switch to get beaten with

OpenStudy (anonymous):

I initially read it as Queenofwelfare, but had to do a double take

OpenStudy (anonymous):

YEP! lol because if you dont know HOW to make a review your screwed

OpenStudy (anonymous):

HAHA! @wio

OpenStudy (anonymous):

I also thought this was going to be about matrices as well, somehow

OpenStudy (anonymous):

4 more problems to go guys! @satellite73